OpenStudy (friendzone):

a number multiplied by -6, subtracted from the sum of 10 and five times the number

OpenStudy (anonymous):

-6x-10 times 5x?

OpenStudy (anonymous):

so what is the question?

OpenStudy (friendzone):

yeah then you gotta convert it you either get a. 11x-10 b. -x+10 c. -x-10 d. 11x+10

OpenStudy (anonymous):

10-6x=5x(is it equal sign there?) 11x-10=0 it semms similar to choice a!!

OpenStudy (anonymous):

I'm thinking a

OpenStudy (anonymous):

I think A same @hayhay5270 But it can also be something different.. these are tricky questions for practically everyone. click a and tell us what you get

OpenStudy (jhannybean):

break it down: a number multiplied by -6, subtracted from the sum of 10 and five times the number "number multiplied by -6 : -6x "subtracted from" : -6x - "the sum of 10 and five times the number " : (10+5x) NOw bring all these components together.

OpenStudy (anonymous):

Yeah I got alittle confused on the sum and subtracted part but I corrected my answer to get a

OpenStudy (anonymous):

@Korey_Minter

OpenStudy (friendzone):

i got it figured it was b

OpenStudy (anonymous):

i think i made a mistake on changing to equation it could be (10+5x)-(-6x)=10+11x-------D

OpenStudy (jhannybean):

It is : -6x - (10+5x)

OpenStudy (anonymous):

no it is subtracted from 10+5x